### v1

@@ 4.2   Benchmarks
We benchmark RL hedging strategies against the performance of three standard models: (i) the practitioner’s Black-Scholes delta hedging (BS) which applies the current implied volatility into the Black-Scholes formula to obtain the hedged option’s delta; (ii) the Leland, (1985) delta (BSL), which modifies the BS delta to reflect proportional transaction costs; and (iii) the Smile-implied delta (SI) introduced by Bates, (2005) and whose closed-form expression for the IV model (5) is provided by François et al., (2022). The explicit formulas for these three benchmarks are detailed in Appendix C.
@@ 4.3.1   Neural network architecture
We employ the RNN-FNN architecture from Section 2.2.1 with two LSTM cells ($L_{1}=2$) of width 56 ($d_{i}=56$ for $i=1,2$) and two FFNN-hidden layers ($L_{2}=2$) of width 56 with ReLU activation function (i.e., $g_{\mathcal{L}_{i}}(x)=\mbox{max}(0,x)$ for $i=1,2$). In the context of the output FFNN layer $\mathcal{L}_{J}$, which maps the output of hidden layers $Z^{(\mathcal{L},\,L_{2})}_{t}$ into the position in the underlying asset $\delta_{t}$, the standard deep hedging framework typically employs a linear activation function. However, this activation function tends to induce RL agents to adopt doubling strategies, increasing the position in the underlying asset several orders of magnitude over the current position on each period after any loss until the cumulative loss amount is fully recovered. Such strategies are definitely undesirable from the perspective of sound risk management. Therefore, we opt to introduce a leverage constraint through the output layer activation function.
@@
[…]
Appendix D.1 provides numerical experiments that illustrate the necessity of including a leverage constraint in the network architecture to alleviate the aforementioned issues.
Our numerical experiments demonstrate the superiority of the RNN-FNN architecture over standalone LSTM and FFNN networks, considering a leverage constraint of $B=100$ for all agents. These results are presented in Appendix D.2. Additionally, agents are trained as described in Section 2.2.2 on a training set of 400,000 independent simulated paths with mini-batch size of 1,000 and a learning rate of 0.0005 that is progressively adapted by the ADAM optimization algorithm. [^7] ADAM is a dynamic learning rate algorithm engineered to accelerate training speeds in deep neural networks and achieve rapid convergence, details can be seen in Goodfellow et al., (2016). In addition, we include a regularization method called dropout with parameter $p=0.5$ to reduce the likelihood of overfitting and enhance the model performance on unseen data. [^8] Full details of the dropout regularization method can be seen in Goodfellow et al., (2016). The selection of $p=0.5$ stems from our numerical experiments detailed in Appendix D.3, indicating that this value consistently outperforms others across all penalty functions. The training procedure is implemented in Python, using Tensorflow and considering the Glorot and Bengio, (2010) random initialization of the initial parameters of the neural network. Numerical results are obtained from a test set of 100,000 independent paths.
@@ 4.3.2   State space selection
However, implementing numerical methods within the deep policy gradient approach across the entire state space may lead to overfitting. In fact, other studies utilizing the RL framework have reported optimal results with a reduced state space. For instance, Buehler et al., (2019) and Cao et al., (2023) employ the deep hedging algorithm without including the portfolio value $V_{t}^{\delta}$ in the state space configuration. Similarly, Carbonneau, (2021) employs the deep hedging framework in a frictionless market without including $\delta_{t}$ in the state space.
We explore three different state space configurations. The first configuration aims to replicate the full state space, denoted by
@@
  • Results are computed using 100,000 out-of-sample paths according to the conditions outlined in Section 4.3.1. Each panel illustrates the optimal penalty function value of an RL agent considering four transaction cost levels. These values are normalized by the estimated values of each penalty function obtained with BSL delta hedging for each transaction cost level. Agents are trained under the specified penalty functions considering three state spaces: full state space $(V_{t}^{\delta},\delta_{t},\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ (red curve), and the two reduced state spaces $(\delta_{t},\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ (green curve) and $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ (blue curve).
Figure 1 illustrates the estimated value of the four penalty functions in proportion to that obtained with the BSL delta, $\rho(\xi_{T}^{\delta})/\rho(\xi_{T}^{BSL})$, across different transaction cost rates for all agents. These numerical results reveal that RL agents outperform BSL delta hedging for all considered state space configurations, as the relative values of the penalty functions are substantially lower than the reference line value of 1.
@@ 4.4   Benchmarking of hedging strategies
In this section, we compare the performance of RL agents with classic hedging approaches. Aggregated results are presented in Section 4.4.1, while the segmentation among various states of the economy is shown in Section 4.4.2. The following subsections break down the performance of the RL agents with respect to moneyness and maturity of the hedged option, or to transaction cost levels, demonstrating the robustness of the RL approach. Finally, Section 4.6 outlines a comparison of performance over historical paths spanning from January 5, 1996, to December 31, 2020.
@@ 4.4.1   Benchmarking over randomized economic conditions
[…]
Table 1: Aggregated hedging metrics for a short position in an ATM call option with maturity of 63 days.
@@
Metric BS SI CVaR95% CVaR99% MSE SMSE
@@
Avg P&L/$\mbox{CVaR}_{\alpha}(\xi_{T}^{\delta}-E[\xi_{T}^{\delta}])$ 0.165 0.166 0.231 0.187 0.161 0.227
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Agents are trained under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ according to the conditions outlined in Section 4.3.1. BS stands for Black-Scholes delta hedging, whereas SI is the smile-implied delta hedging.
Table 1 presents hedging performance metrics attained by both benchmarks and RL agents. Every RL agent achieves the lowest value for the corresponding metric which they used as objective function during training, which is expected. Furthermore, the numerical results indicate that RL agents provide hedging strategies that are much less risky than benchmarks, as evidenced by metrics such as CVaR95%, CVaR99%, and SMSE. In particular, when computing the variation rate between the estimated value of each of these metrics obtained by RL agents relative to the value obtained by BS delta, we observe an average reduction rate across RL agents of 24%, 31%, and 38%, respectively. Similarly, when comparing with SI delta strategies, we observe average reductions across RL agents of 45%, 46%, and 67%, respectively. Moreover, in terms of the MSE metric, the MSE agent demonstrates significant superiority over benchmarks, reducing 26% and 63% compared to BS delta and SI delta strategies, respectively.
Regarding average Profit and Loss (P&L), SI delta yields higher profitability; however, it also entails increased risk. In contrast, RL agents trained using CVaR95%, CVaR99%, and SMSE as penalty functions obtain lower values compared to SI delta strategies, but they reach a more favorable trade-off between profitability and risk management as shown by the ratio Avg P&L/$\mbox{CVaR}_{\alpha}(\xi_{T}^{\delta}-E[\xi_{T}^{\delta}])$, which yields greater values for these agents. Additionally, RL agents achieve lower risk than both the BS and SI delta hedging strategies, which highlights enhanced risk management.
@@ 4.4.2   Impact of the state of the economy on performance
Table 2: Clusters of dates representing different time periods
@@
Crisis 2 (Dot-com bubble crisis) 03/$0$ 31/12/2002 \$5.758
@@
Our numerical results, illustrated in Figure 2, indicate that both benchmarks and RL agents are sensitive to the environment, showcasing better performance during regular periods compared to financial crises in terms of risk, as shown by the two right plots showing the CVaR95% and SMSE metrics. Additionally, both approaches tend to offer more profitable strategies during crisis periods and periods of high volatility (for instance, even though Regular period 2 is not labeled as a crisis periods, it is characterized by economic recovery with significant market fluctuations). This trend can be explained by the fact that higher volatility during these periods leads to a higher risk premium, increasing the initial portfolio value and thus the final P&L, as highlighted by the Avg P&L metric (see the top left panel of Figure 2). Regarding the MSE metric, the results align with the Avg P&L metric, with both approaches showing superior performance during regular periods (except for Regular period 2 which exhibited high volatility), as higher profits are detrimental for the MSE metric penalizing upside risk.
@@
The results that are segregated by time period align with the aggregated findings. Indeed, for any of the CVaR95%, MSE, and SMSE risk metrics, the RL agent which was trained with such risk metric as its objective function consistently exhibits lower values than benchmarks across all periods. These differences are particularly notable during crisis periods, where BS delta and SI delta tend to be riskier, while RL agents demonstrate greater stability. The higher values in the MSE statistic of agents trained under CVaR and SMSE can be attributed to the fact that these objective functions only penalize hedging losses, allowing agents to seek positive returns on average. This is consistent with the Avg P&L panel, where CVaR and SMSE agents tend to display more profitable strategies than the MSE agent.
@@ 4.4.3   Impact of moneyness level on performance
In line with our previous experiments, we observe that SI delta tends to offer more profitable strategies compared to benchmarks, while BS delta demonstrates a similar level of profitability for ATM and ITM options. However, this trend does not hold for OTM options, where RL agents not only exhibit better risk management but also yield more profitable strategies on average (refer to CVaR and SMSE agents for OTM options in Figure 3). The discrepancy can be attributed to the fact that RL agents trained under CVaR and SMSE do not track the option value nor penalize gains at maturity, allowing agents to profit from OTM paths at maturity. In contrast, BS and BSL are option tracking methodologies that aim to replicate the option value at maturity, thereby reducing potential gains for OTM options at maturity. The latter is consistent with the RL agent trained under the MSE, which achieves the minimum MSE value for OTM options and the lowest Avg P&L, as it penalizes both gains and losses.
@@ 4.4.5   Impact of transaction costs on performance
We now investigate scenarios where hedgers encounter transaction costs, which is meant to reflect more realistic hedging scenarios. Specifically, we analyze the hedging effectiveness of RL agents in comparison to benchmarks, including the BSL delta hedging. Our examination focuses on hedging an ATM European call option with a maturity of $N=63$ days across varying levels of transaction costs assumed to be 0.05%, 0.5% and 1%.
Figure 5 illustrates the performance of benchmarks and RL agents, with the four panels depicting the Avg P&L, CVaR95%, MSE, and SMSE metrics. In general, as expected, the performance of each strategy tends to deteriorate as transaction costs increase. However, performance drops are more pronounced for benchmarks than for RL agents. As anticipated, RL agents consistently outperform benchmarks in terms of downside risk (observe CVaR95% and SMSE metrics depicted in the panels positioned at the top and bottom right corners of Figure 5) across all transaction cost levels, regardless of the penalty function used during training. Furthermore, while BSL delta hedging adjusts the strategies in terms of the transaction cost level and performs better than the other benchmarks, RL strategies display superior metrics, highlighting their better capacity to adapt to different transaction cost levels.
@@
  • Results are computed using 100,000 out-of-sample paths under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ without transaction costs, and under $(\delta_{t},\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ in the presence of transaction cost. Agents are trained according to the conditions outlined in Section 4.3.1. The average option price is \$3.89 with a standard deviation of \$1.29. The acronym TC stands for transaction cost. BS stands for Black-Scholes delta hedging, BSL for eland delta hedging and SI for smile-implied delta hedging.
@@
across different transaction cost levels. In the cases of BS delta and SI delta strategies, which overlook transaction costs, we observe a higher cost of hedging due to significant variations in the position of the underlying asset during rebalancing across different time steps, thereby leading to lower profitability as transaction costs increase. This trend is evident in the case of SI delta which incurs the highest average hedging cost, followed by BS delta and BSL delta, when the transaction cost rate is set to 0.5% and 1%. These results consistently align with the performance differences observed in Figure 5, where benchmarks are associated with the highest losses.
@@
$\kappa$ Metric BS SI Leland CVaR95% CVaR99% MSE SMSE
@@
4.5   Global importance of IV surface features
We now investigate to what extent the risk factors characterizing the IV surface described in Equation (5) contribute to total performance of our model. We employ the Shapley Additive Global Importance (SAGE) methodology, introduced by Covert et al., (2020), to evaluate their impact. Specifically, let $\rho\left(\xi_{T}^{\tilde{\delta}_{\theta}(\cal{X})}\right)$ be the risk measure when the model is trained with the state space $\cal{X}$. The amount of risk reduction achieved by adding the state variables ($(\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ to a baseline model $\tilde{\delta}_{\theta}$ with state variables $(\tau_{t},S_{t})$ is

### v2

@@ 4.2  Benchmarks
We benchmark RL hedging strategies against the performance of three delta hedging strategies using the following deltas: (i) the practitioner’s Black-Scholes delta hedging (DH) which applies the current implied volatility into the Black-Scholes formula to obtain the hedged option’s delta; (ii) the Leland, (1985) delta (DH-L), which modifies the DH to reflect proportional transaction costs; and (iii) the Smile-implied delta (SI) introduced by Bates, (2005) and whose closed-form expression for the IV model (5) is provided by François et al., (2022). The explicit formulas for these three benchmarks are detailed in Appendix A.
@@ 4.3.1 Neural network architecture
We employ the RNN-FNN architecture from Section 2.2.1 with two LSTM cells ($L_{1}=2$) of width 56 ($d_{i}=56$ for $i=1,2$) and two FFNN-hidden layers ($L_{2}=2$) of width 56 with ReLU activation function (i.e., $g_{\mathcal{L}_{i}}(x)=\mbox{max}(0,x)$ for $i=1,2$). In the context of the output FFNN layer $\mathcal{L}_{J}$, which maps the output of hidden layers $Z^{(\mathcal{L},\,L_{2})}_{t}$ into the position in the underlying asset $\delta_{t}$, the standard deep hedging framework typically employs a linear activation function. However, this activation function tends to induce RL agents to adopt doubling strategies, increasing the position in the underlying asset several orders of magnitude over the current position on each period after any loss until the cumulative loss amount is fully recovered. Such strategies are definitely undesirable from the perspective of sound risk management. Therefore, we opt to introduce a leverage constraint through the output layer activation function. This leverage constraint is a dynamic upper bound on the activation function, denoted as $g_{\mathcal{L}_{J}}(Z,t)=\mbox{min}(Z,B_{t}(Z))$, with $Z=T_{\mathcal{L}_{J}}(Z^{(\mathcal{L},\,L_{2})}_{t})$ representing the typical deep hedging underlying asset position, and $B_{t}(Z)$ the highest position in the index that can be held in the portfolio. Such an upper bound avoids excessive leverage and limits the borrowing capacity, i.e., the cash held satisfies $\phi_{t+1}(X_{t})\geq-B$ for all $X_{t}$ and $t$, and B>0B>0.888This type of leverage condition has been previously employed in the literature. For instance, the seminal paper of Harrison and Pliska, (1981) assumes a restricted borrowing capacity to maintain a positive wealth throughout the entire hedging period. The latter, in conjunction with the self-financing constraint (1), establishes the dynamic upper bound as
@@
Appendix E.1 of the supplementary material provides numerical experiments that illustrate the necessity of including a leverage constraint in the network architecture to alleviate the aforementioned issues.
Our numerical experiments demonstrate the superiority of the RNN-FNN architecture over standalone LSTM and FFNN networks, considering a leverage constraint of $B=100$ for all agents. These results are presented in Appendix E.2 of the supplementary material. Additionally, agents are trained as described in Section 2.2.2 on a training set of 400,000 independent simulated paths with mini-batch size of 1,000 and a learning rate of 0.0005 that is progressively adapted by the ADAM optimization algorithm. [^9] ADAM is a dynamic learning rate algorithm engineered to accelerate training speeds in deep neural networks and achieve rapid convergence, details can be seen in Goodfellow et al., (2016). In addition, we include a regularization method called dropout with parameter $p=0.5$ to reduce the likelihood of overfitting and enhance the model performance on unseen data. [^10] Full details of the dropout regularization method can be seen in Goodfellow et al., (2016). The selection of $p=0.5$ stems from our numerical experiments detailed in Appendix E.3 of the supplementary material, indicating that this value consistently outperforms others across all penalty functions. The training procedure is implemented in Python, using Tensorflow and considering the Glorot and Bengio, (2010) random initialization of the initial parameters of the neural network. Numerical results are obtained from a test set of 100,000 independent paths.
@@ 4.3.2 State space selection
However, implementing numerical methods within the deep policy gradient approach across the entire state space may lead to overfitting. In fact, other studies utilizing the RL framework have reported optimal results with a reduced state space. For instance, Buehler et al., (2019) and Cao et al., (2023) employ the deep hedging algorithm without including the portfolio value $V_{t}^{\delta}$ in the state space configuration. Similarly, Carbonneau, (2021) employs the deep hedging framework in a frictionless market without including $\delta_{t}$ in the state space. We explore three different state space configurations. The first configuration aims to replicate the full state space, denoted by
@@
  • Results are computed using 100,000 out-of-sample paths according to the conditions outlined in Section 4.3.1. Each panel illustrates the optimal penalty function value of an RL agent considering four transaction cost levels. These values are normalized by the estimated values of each penalty function obtained with DH-L for each transaction cost level. Agents are trained under the specified penalty functions considering three state spaces: full state space $(V_{t}^{\delta},\delta_{t},\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ (red curve), and the two reduced state spaces $(\delta_{t},\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ (green curve) and $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ (blue curve).
Figure 1 illustrates the estimated value of the four penalty functions in proportion to that obtained with DH-L, $\rho(\xi_{T}^{\delta})/\rho(\xi_{T}^{\text{DH-L}})$, across different transaction cost rates for all agents. These numerical results reveal that RL agents outperform DH-L for all considered state space configurations, as the relative values of the penalty functions are substantially lower than the reference line value of 1.
@@ 4.4  Benchmarking of hedging strategies
In this section, we compare the performance of RL agents with classic hedging approaches. Aggregated results are presented in Section 4.4.1, while the segmentation among various states of the economy is shown in Section 4.4.2. The following subsections break down the performance of the RL agents with respect to moneyness and maturity of the hedged option, or to transaction cost levels, demonstrating the robustness of the RL approach. Finally, Section 5 outlines a comparison of performance over historical paths spanning from January 5, 1996, to December 31, 2020.
@@ 4.4.1 Benchmarking over randomized economic conditions
We begin by comparing the hedging performance of the benchmarks and RL agents trained under the four penalty functions: CVaR95%, CVaR99%, MSE, and SMSE. This comparison is performed in terms of estimated values of all penalty functions and the average Profit and Loss, $\mbox{Avg P\&L}=\mathbb{E}[-\xi_{T}^{\delta}]$, across all paths in a test set. Additionally, we employ the CVaR deviation measure, defined as $\mbox{CVaR}_{\alpha}(\xi_{T}^{\delta}-E[\xi_{T}^{\delta}])$, as a deviation metric for the hedging error in the test set. In such test set, initial economic conditions (the initial value of state variables in the path) are sampled randomly among historical values from our sample. It therefore reflects aggregate performance across various economic conditions. Our analysis focuses on hedging a European ATM call option with a maturity of $N=63$, under the assumption of no transaction costs, which is $\kappa=0$.
Table 1: Aggregated hedging metrics for a short position in a 63-day ATM call option.
@@
Metric DH SI CVaR95% CVaR99% MSE SMSE
@@
[…]
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Agents are trained under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ according to the conditions outlined in Section 4.3.1. DH stands for delta hedging, whereas SI is the smile-implied delta hedging.
Table 1 presents hedging performance metrics attained by both benchmarks and RL agents. Every RL agent achieves the lowest value for the corresponding metric which they used as objective function during training, which is expected. Furthermore, the numerical results indicate that RL agents provide hedging strategies that are much less risky than benchmarks, as evidenced by metrics such as CVaR95%, CVaR99%, and SMSE. In particular, when computing the variation rate between the estimated value of each of these metrics obtained by RL agents relative to the value obtained by DH, we observe an average reduction rate across RL agents of 24%, 31%, and 38%, respectively. Similarly, when comparing with SI delta strategies, we observe average reductions across RL agents of 45%, 46%, and 67%, respectively. Moreover, in terms of the MSE metric, the MSE agent demonstrates significant superiority over benchmarks, reducing 26% and 63% compared to DH and SI delta strategies, respectively.
[…]
@@ 4.4.2 Impact of the state of the economy on performance
Table 2: Clusters of dates representing different time periods.
@@
Crisis 2 (Dot-com bubble crisis) 03/01/2000 - 31/12/2002 \$5.758
@@
Our numerical results, illustrated in Figure 2, indicate that both benchmarks and RL agents are sensitive to the environment, showcasing better performance during regular periods compared to financial crises in terms of risk, as shown by the two right plots showing the CVaR95% and SMSE metrics. Additionally, both approaches tend to offer more profitable strategies during crisis periods and periods of high volatility (for instance, even though Regular period 2 is not labeled as a crisis period, it is characterized by economic recovery with significant market fluctuations). This trend can be explained by the fact that higher volatility during these periods leads to a higher risk premium, increasing the initial portfolio value and thus the final P&L, as highlighted by the Avg P&L metric (see the top left panel of Figure 2). Regarding the MSE metric, the results align with the Avg P&L metric, with both approaches showing superior performance during regular periods (except for Regular period 2 which exhibited high volatility), as higher profits are detrimental for the MSE metric penalizing upside risk.
@@
The results that are segregated by time period align with the aggregated findings. Indeed, for any of the CVaR95%, MSE, and SMSE risk metrics, the RL agent which was trained with such risk metric as its objective function consistently exhibits lower values than benchmarks across all periods. These differences are particularly notable during crisis periods, where DH and SI delta tend to be riskier, while RL agents demonstrate greater stability. The higher values in the MSE statistic of agents trained under CVaR and SMSE can be attributed to the fact that these objective functions only penalize hedging losses, allowing agents to seek positive returns on average. This is consistent with the Avg P&L panel, where CVaR and SMSE agents tend to display more profitable strategies than the MSE agent.
@@ 4.4.3 Impact of moneyness level on performance
In line with our previous experiments, we observe that SI delta tends to offer more profitable strategies compared to benchmarks, while DH demonstrates a similar level of profitability for ATM and ITM options. However, this trend does not hold for OTM options, where RL agents not only exhibit better risk management but also yield more profitable strategies on average (refer to CVaR and SMSE agents for OTM options in Figure 3). The discrepancy can be attributed to the fact that RL agents trained under CVaR and SMSE do not track the option value nor penalize gains at maturity, allowing agents to profit from OTM paths at maturity. In contrast, DH and SI are option tracking methodologies that aim to replicate the option value at maturity, thereby reducing potential gains for OTM options at maturity. The latter is consistent with the RL agent trained under the MSE, which achieves the minimum MSE value for OTM options and the lowest Avg P&L, as it penalizes both gains and losses.
@@ 4.4.5 Impact of transaction costs on performance
We now investigate scenarios where hedgers encounter transaction costs, which is meant to reflect more realistic hedging scenarios. Specifically, we analyze the hedging effectiveness of RL agents in comparison to benchmarks, including the DH-L strategy. Our examination focuses on hedging an ATM European call option with a maturity of $N=63$ days across varying levels of transaction costs assumed to be 0.05%, 0.5% and 1%.
Figure 5 illustrates the performance of benchmarks and RL agents, with the four panels depicting the Avg P&L, CVaR95%, MSE, and SMSE metrics. In general, as expected, the performance of each strategy tends to deteriorate as transaction costs increase. However, performance drops are more pronounced for benchmarks than for RL agents. As anticipated, RL agents consistently outperform benchmarks in terms of downside risk (observe CVaR95% and SMSE metrics depicted in the panels positioned at the top and bottom right corners of Figure 5) across all transaction cost levels, regardless of the penalty function used during training. Furthermore, while DH-L adjusts the strategies in terms of the transaction cost level and performs better than the other benchmarks, RL strategies display superior metrics, highlighting their better capacity to adapt to different transaction cost levels.
@@
  • Results are computed using 100,000 out-of-sample paths under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ without transaction costs, and under $(\delta_{t},\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ in the presence of transaction cost. Agents are trained according to the conditions outlined in Section 4.3.1. The average option price is \$3.89 with a standard deviation of \$1.29. The acronym TC stands for transaction cost. DH Black-Scholes delta hedging, DH-L for Leland delta hedging and SI for smile-implied delta hedging.
@@
across different transaction cost levels. In the cases of DH and SI delta strategies, which overlook transaction costs, we observe a higher cost of hedging due to significant variations in the position of the underlying asset during rebalancing across different time steps, thereby leading to lower profitability as transaction costs increase. This trend is evident in the case of SI delta which incurs the highest average hedging cost, followed by DH and DH-L, when the transaction cost rate is set to 0.5% and 1%. These results consistently align with the performance differences observed in Figure 5, where benchmarks are associated with the highest losses.
@@
$\kappa$ Metric DH SI DH-L CVaR95% CVaR99% MSE SMSE
@@
4.5  Global importance of IV surface factors
We now investigate to what extent the risk factors characterizing the IV surface described in Equation (5) contribute to total performance of our model. We employ the Shapley Additive Global Importance (SAGE) methodology, introduced by Covert et al., (2020), to evaluate their impact. Specifically, let $\rho\left(\xi_{T}^{\tilde{\delta}_{\theta}(\cal{X})}\right)$ be the risk measure when the model is trained with the state space $\cal{X}$. The amount of risk reduction achieved by adding the state variables $(\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ to a baseline model $\tilde{\delta}_{\theta}$ with state variables $(\tau_{t},S_{t})$ is